Find an equation of the line with slope 3 and containing the point of intersection of 2x+3y-5=0 and 3x-7y+5=0.
Answer by lwsshak3(11628)

Find an equation of the line with slope 3 and containing the point of intersection of 2x+3y-5=0 and 3x-7y+5=0.
**
3x-7y+5=0
2x+3y-5=0
..
6x-14y+10=0 (multiply by 2)
6x+9y-15=0 (multiply by 3)
subtract
-23y+25=0
y=25/23≈1.087
2x=5-3y=5-75/23
x=(5-75/23)/2≈0.870
point of intersection: (0.870,1.087)
equation of line: y=mx+b
y=3x+b
solve for b using coordinates of point of intersection
1.087=3*.870+b
b=-1.523
y=3x-1.523
comment: what horrible answers!